OpenStudy (anonymous):

evaluate the expressions as rational numbers by using properties of exponents and logs? ln(1/(e^2))

OpenStudy (anonymous):

the first rule to use would be the difference rule ln1-lne^2, then cross out the ln and e because they are inverses of each other ln1-2

OpenStudy (luffingsails):

Christal, The key to log is understanding some of the basics. Consider some of the following: \[\ln {ab} = \ln{a} + \ln{b}\] \[\ln{a/b} = \ln{a}-\ln{b}\] \[\ln{e^n} = n\] So, just re-write the original equation using those laws: \[\ln(1/e^2) = \ln{1}-\ln{e^2}\] From the second and third rules above, it simplifies to: \[\ln{1}-2\] Hope that helps

OpenStudy (anonymous):

it did! thank you